OpenStudy (agentnao):

Calc 2 Help: Derivative of Inverse Trig Functions y = xarcsin(x/4) + sqrt(16-x^2)

OpenStudy (mathmale):

Dear Agent-N: Appreciate the way you use parentheses appropriately to ensure proper interpretation of these math expressions. It appears to me that your xarcsin(x/4) is a product. Correct?

OpenStudy (agentnao):

Yes! That is correct.

OpenStudy (mathmale):

Do you have available a reference work that lists the derivatives of inverse trig functions?

OpenStudy (agentnao):

I know the derivatives of arcsin, arccos, and arctan.

OpenStudy (mathmale):

The derivative with respect to x of arcsin(x) is 1 over Sqrt(1-x^2). Agreed?

OpenStudy (usukidoll):

don't we need the product rule as well for xarcsin(x/4)?

OpenStudy (usukidoll):

like f g' + g f'?

OpenStudy (agentnao):

Oh. I forgot to include the rest of the question. That is the function, but what they want me to do is find y'(2)

OpenStudy (usukidoll):

o-o

OpenStudy (usukidoll):

oh I see we need the derivative first and then plug in y'(2)

OpenStudy (mathmale):

Yes, of course; but the focus of this problem is "inverse trig functions," and so I'm reviewing the derivative of arcsin x before applying the product rule.

OpenStudy (agentnao):

I used the product rule, but I still have a arcsin in the derivative that I don't know how to handle.

OpenStudy (usukidoll):

it should be in your textbook. they're given in a chart.

OpenStudy (usukidoll):

http://upload.wikimedia.org/math/9/a/6/9a6ca82be93f7fd7a2b8d102a308b547.png use the first one

OpenStudy (mathmale):

That makes complete sense, AgentN, that you still have the arcsin function in your expression. Simply leave it as is there; you are writing the derivative of the arcsin function in another term, right?

OpenStudy (agentnao):

I know the derivatives of inverse trig functions. That is not the problem.

OpenStudy (mathmale):

Would you mind defining what IS posing a problem for you?

OpenStudy (agentnao):

Not in another term. I need to calculate the value of the expression at 2. After finding the derivative, of course.

OpenStudy (usukidoll):

I guess the arcsin is left alone.. like we leave the x alone deal with the arcsin + leave the arcsin alone deal with the x

OpenStudy (usukidoll):

and then plug in ... oh I see how to calculate arcsin

OpenStudy (agentnao):

Here. Let me tell you what I have right now.

OpenStudy (usukidoll):

calculating the arcsin part when you put y'(2)....

OpenStudy (mathmale):

Good approach. It'd help me tremendously to see what you have done so far.

OpenStudy (agentnao):

x/[4sqrt(1-x^2)] + arcsin(x/4) - 1/sqrt(16-x^2)

OpenStudy (agentnao):

That is the derivative I came up with, but I don't know how to plug in 2 to arcsin.

OpenStudy (usukidoll):

2/[4sqrt(1-2^2)] + arcsin(2/4) - 1/sqrt(16-2^2)

OpenStudy (usukidoll):

2/[4sqrt(1-4)] + arcsin(1/2) - 1/sqrt(12)

OpenStudy (agent0smith):

arcsin(2/4) is fine... arcsin2 is not.

OpenStudy (mathmale):

I believe you are completely on the right track there. Now you just want to evaluate this expression at x=2, right?

OpenStudy (agent0smith):

Oh, wait.

OpenStudy (usukidoll):

2/[4sqrt(-3)] + arcsin(1/2) - 1/sqrt(12)

OpenStudy (agentnao):

Yes. And my other problem is the negative sign under the radical.

OpenStudy (agent0smith):

Yeah, arcsin(1/2) is fine.... the square root, this just means the derivative doesn't exist at x=2.

OpenStudy (usukidoll):

derivative doesn't exist at x = 2... because of that negative sign?

OpenStudy (agent0smith):

Assuming the derivative is calculated correctly...

OpenStudy (mathmale):

Let's focus on the very first term. x/[4sqrt(1-x^2)] + arcsin(x/4) - 1/sqrt(16-x^2) Replacing x with 2 results in a first term that looks like 2 ---------------------- 2*Sqrt(1-(2/4)^2) Think: where did that (2/4) come from, Agent N? Yes, assuming that the derivative was calculated correctly, agentOsmith..

OpenStudy (mathmale):

Agent N: Please quickly review your original statement of this homework problem and note that the argument of arcsin is (x/4), not just x.

OpenStudy (agent0smith):

^i don't think it's correct. derivative of arcsin(x/4) = 1/sqrt ( 1-(x/4)^2 )

OpenStudy (agentnao):

Definitely x/4!

OpenStudy (mathmale):

Thus, if x=2, we end up with arcsin (2/4) = arcsin (1/2) = pi/6 radians.

OpenStudy (agentnao):

Are you sure, agent0smith?? I thought about that, but inserting x/4 as the input to the inverse function would make it a chain rule. Right?

OpenStudy (usukidoll):

uhh isn't sin 30 = pi/6 that's different than arcsin

OpenStudy (agent0smith):

Oh yeah i forgot about that. 1/4*1/sqrt ( 1-(x/4)^2 )

OpenStudy (agent0smith):

=1/sqrt(16 - x^2)

OpenStudy (mathmale):

Once we have a tentative solution, I'd suggest that you, Agent N, quickly re-do the problem to ensure that all the differentiation (including application of the chain rule) has been done correctly.

OpenStudy (agent0smith):

Check it here: http://www.wolframalpha.com/input/?i=arcsin%28x%2F4%29++derivative

OpenStudy (agentnao):

Do you think we'll figure it out within 20 minutes? Because that's when this question is due :(

OpenStudy (usukidoll):

:O

OpenStudy (agent0smith):

^yes. we have already.

OpenStudy (mathmale):

arcsin (1/20 = Pi/6 (an angle) sin (Pi/6) = sin (30 deg) = 1/2 (a ratio).

OpenStudy (agentnao):

I see agentsmith. I completely forgot about good ole wolfphram.

OpenStudy (agentnao):

So what is the final answer then?

OpenStudy (mathmale):

OK, Agent N: seeing that you have 3 helpers at once, perhaps it'd be good for you to sit back for a moment and tell us exactly what it is that we can help you to finish this problem solution.

OpenStudy (agentnao):

Is my derivative correct? WIth the negative under the radical symbol?

OpenStudy (agent0smith):

Derivative of: y = xarcsin(x/4) + sqrt(16-x^2) y' = arcsin(x/4) +x*1/sqrt(16 - x^2) + 0.5*2x/sqrt(16-x^2) w/o really simplifying. Then plug in x=2

OpenStudy (agentnao):

And the 2/4 as the arcsin input?

OpenStudy (agent0smith):

There is no negative sign

OpenStudy (mathmale):

The 2/4 as input to the arcsin function is just fine. I think you're referring to the derivative of arcsin (x-4), which DOES have a negative sign under the radical.

OpenStudy (agent0smith):

or alternatively: http://www.wolframalpha.com/input/?i=Derivative+of%3A+y+%3D+xarcsin%28x%2F4%29+%2B+sqrt%2816-x%5E2%29 now plug in x=2

OpenStudy (agentnao):

I see. I had the wrong derivative of arcsin(x/4). That takes care of it. I think. Let me calculate it fully now.

OpenStudy (mathmale):

Excuse me: you've probably noticed I've made a typo or two. arcsin (x/4) is correct.

OpenStudy (mathmale):

Yes, why don't you take your time and then ask questions about anything remaining unclear to you.

OpenStudy (agentnao):

Yes. I'll be back in just one second.

OpenStudy (agentnao):

Shoot. I'm not getting it.

OpenStudy (usukidoll):

I wish the draw function would come back T_T... can you scan what you have and attach it on here?

OpenStudy (mathmale):

As before, please pinpoint where you need help.

OpenStudy (agentnao):

I plugged in 2 for the wolfphram alpha derivative and I got a string of radical expressions and pi/6 that wasn't correct :(

OpenStudy (agentnao):

My arithmetic might be wrong.

OpenStudy (mathmale):

What can I/we do to help you? Please take charge.

OpenStudy (inkyvoyd):

They want y'(2), correct?

OpenStudy (usukidoll):

k k kk k k the best thing to do right now. is to take the derivative.. product rule slowly.... for the first one.

OpenStudy (mathmale):

Yes: first find the derivative of the given function, and then evaluate that derivative at x=2.

OpenStudy (agentnao):

Yes. What is the final value.

OpenStudy (agentnao):

I have 5 minutes left.

OpenStudy (usukidoll):

:S the pressure... ok ok .. errr

OpenStudy (inkyvoyd):

how many tries?

OpenStudy (agentnao):

Im usually all for taking my time, but I spent an hour on this.. I will make sure to understand it later but I need help

OpenStudy (agent0smith):

arcsin(1/2) = ? if theta = arcsin(1/2) then sin theta = 1/2 so theta = 30 degrees or pi/6

OpenStudy (agentnao):

inkyvoyd! 3 tries left.

OpenStudy (usukidoll):

I think it's one of those timed assignments like mymathlab

OpenStudy (agentnao):

It's webassign. The due date for the assignment is midnight.

OpenStudy (agentnao):

I think I'll have to give this one up.

OpenStudy (agent0smith):

I just wrote the solution!

OpenStudy (inkyvoyd):

screenshot and crop the problem?

OpenStudy (agentnao):

The final answer is not pi/6. What about the rest of it?

OpenStudy (usukidoll):

I KNOW! @dan815 SAVE US D:

OpenStudy (agent0smith):

What rest? It is pi/6.

OpenStudy (agentnao):

http://screencast.com/t/NPK40sA4ltE4

OpenStudy (inkyvoyd):

and you put down just pi/6 and it was wrong?

OpenStudy (agent0smith):

THE FINAL ANSWER IS pi/6 AND NOTHING ELSE

OpenStudy (agentnao):

Oh....you were right agent smith. I'm sorry...

OpenStudy (agent0smith):

lol i gave you this: or alternatively: http://www.wolframalpha.com/input/?i=Derivative+of%3A+y+%3D+xarcsin%28x%2F4%29+%2B+sqrt%2816-x%5E2%29 now plug in x=2

OpenStudy (agentnao):

THANK YOU GUYS SO MUCH. I'm still not sure WHY it's pi/6, but I have other obligations for the evening, so I will look into it tomorrow morning. Thanks again for all your help!!

OpenStudy (agent0smith):

y=xarcsin(x/4) + sqrt(16-x^2) I see why. We forgot a negative. y' = arcsin(x/4) + 1/4*x/sqrt(1-(x/4)^2) + 0.5*(-2x)/sqrt(16-x^2) (those last two terms are the same but diff signs, so they cancel)... y' = arcsin(x/4)

OpenStudy (mathmale):

Thank you very much for your extra effort, AOS.

OpenStudy (agent0smith):

^i only noticed when checking the solution on W.A for how they simplified it... i wondered why they had a negative sign. Derivative of -x^2 is -2x... jeez.